OpenStudy (anonymous):

Find, correct to the nearest degree, the three angles of the triangle with the given vertices. A(0, 1, 2), B(-1, 5, 1), C(3, 2, -1),

OpenStudy (anonymous):

calculate AB, and BC vectors, then use formula finding out angle between them to find the answer, it is theta = \[\theta=\cos^- \frac{ AB \dot BC }{ ||AB|| ||BC|| }\] do the same to find out other angles

OpenStudy (anonymous):

I'm not sure how to find those values...